# 2.5D Transformer: An Efficient 3D Seismic Interpolation Method without Full 3D Training

### v1

@@
Abstract
Transformer has emerged as a powerful deep-learning technique for two-dimensional (2D) seismic data interpolation, owing to its global modeling ability. However, its core operation introduces heavy computational burden due to the quadratic complexity, hindering its further application to higher-dimensional data. To achieve Transformer-based three-dimensional (3D) seismic interpolation, we propose a 2.5-dimensional Transformer network (T-2.5D) that adopts a cross-dimensional transfer learning (TL) strategy, so as to adapt the 2D Transformer encoders to 3D seismic data. The proposed T-2.5D is mainly composed of 2D Transformer encoders and 3D seismic dimension adapters (SDAs). Each 3D SDA is placed before a Transformer encoder to learn spatial correlation information across seismic lines. The proposed cross-dimensional TL strategy comprises two stages: 2D pre-training and 3D fine-tuning. In the first stage, we optimize the 2D Transformer encoders using a large amount of 2D data patches. In the second stage, we freeze the 2D Transformer encoders and fine-tune the 3D SDAs using limited 3D data volumes. Extensive experiments on multiple datasets are conducted to assess the effectiveness and efficiency of T-2.5D. Experimental results demonstrate that the proposed method achieves comparable performance to that of full 3D Transformer at a significantly low cost.
@@ Abstract
Seismic data interpolation, Cross-dimensional transfer learning, Transformer, Seismic dimension adapter1 Introduction
@@
Interpolation operation is a crucial step for improving the quality of seismic data, especially when facing geometries with large spatial sampling intervals, and recovering missing parts caused by topographical constraints, limited budgets, or receiver malfunctions. By reconstructing the missing information and compensating the sparse geometries, effective interpolations can enhance the completeness and continuity of events, facilitating the subsequent inversion and high-resolution imaging [chen2024low, cheng2025generative].
@@
To acquire the desired dense and complete seismic data, geophysical scholars have proposed various conventional interpolation methods over the past few decades [spitz1991seismic, trad2002accurate, oropeza2011simultaneous]. These methods are theory-driven, relying on mathematical models or analytical formulations to describe the underlying physical processes. Generally, they can be divided into four categories: the wave-equation-based, prediction filtering, low-rank, and sparse-representation methods.
@@
The wave-equation-based methods utilize subsurface velocity models to simulate the propagation of seismic wavefields, so as to reconstruct the incomplete data [wang2019deep]. [ronen1987wave] proposed a trace-interpolation method based on the wave equation and prior assumption of a smooth spatial spectrum, and experimental results have demonstrated its effectiveness on both synthetic and field data. The offset-continuation differential equation-based method proposed by [fomel2003seismic] showed good interpolation performance even in several structurally complex situations. However, the wave-equation based methods often require a relatively accurate velocity model as an important prerequisite, which is often unavailable in numerous real-world problems. Moreover, the heavy computation burden seriously limits their further applications.
@@
The prediction filtering methods exploit the correlation of events in different domains to perform the interpolation operation. When applying these methods, the interpolation task is taken as a least-squares linear inverse problem, and interpolated results are generated by minimizing the misfit between the predicted and input data [liu2022bseismic]. [spitz1991seismic] proposed a multichannel and model-free interpolation method for missing traces in f-x domain, exhibiting good interpolation performance in both pre- and post-stack examples. To avoid solving the complex linear system of equations in Spitz method, [porsani1999seismic] introduced a half-step predictive filter to make the interpolation operation more efficient and easier for implementation. [wang2002seismic] extended the f-x domain interpolation to f-x-y domain by masking the data samples selectively from the input frequency slice to design the linear prediction (LP) operators. To further enhance the interpolation performance of predictive filters on irregularly sampled data, [naghizadeh2007multistep] proposed a multistep autoregressive algorithm-based prediction-error method combined with Fourier-based methods to reconstruct seismic data from low to all frequencies. An important assumption of these prediction filtering methods is that events are required to be linear. However, this is usually not satisfied in real conditions, resulting in serious interpolation performance degradation.
@@
The low-rank methods assume that the complete seismic data in a specific arrangement is of low rank at a given frequency component [trickett2010, gao2013fast], and missing traces will significantly increase its rank. Therefore, the interpolation tasks can be accomplished by reducing the rank to an optimal one. Typical low-rank methods include singular spectrum analysis [SSA; vautard1989singular, oropeza2011simultaneous, lari2019adaptive, carozzi2021interpolated, li2024interpolated], matrix completion [ma2013three, yang2013seismic, kumar2015efficient], principal component analysis [WOLD198737, li2015inverse, huang2016improved, wu2023high], and Cadzow filtering [cadzow2002signal, gao2013fast, naghizadeh2013multidimensional, huang2020dealiased]. [oropeza2011simultaneous] proposed a simultaneous denoising and reconstruction method for seismic data based on multicahnnel SSA (MSSA), which resembles seismic data interpolation with the method of projection onto convex sets (POCS). Moreover, a randomized singular value decomposition is adopted to accelerate its rank reduction stage. [ma2013three] enhanced the matrix completion using a designed texture-patch transformation, exhibiting superior performance to traditional POCS method. [naghizadeh2013multidimensional] used the low-rank approximation of the Hankel matrix of seismic data to recover missing spatial samples. Its effectiveness was validated in the reconstruction of synthetic and real data examples. However, in these low-rank methods, the determination of optimal rank is an intricate problem, and an improper one will have a negative impact on the interpolated results [ma2013three, cheng2023seismic].
@@
The sparse-representation methods interpolate seismic records by representing them into sparse domains. The useful signals and the missing data are treated as large- and small-amplitude coefficients, respectively. Preserving the large-amplitude coefficients and eliminating the small-amplitude ones helps to extract the useful signal components from the incomplete data [chen2019interpolation]. The missing signal components are then recovered through an inverse sparse transform. These methods can be mainly categorized into mathematical transform-based and dictionary learning-based methods. [trad2002accurate] proposed a high-resolution time-variant Radon transform (RT)-based interpolation method. Hyperbolic and elliptical RTs are implemented to perform accurate interpolation and attenuate sampling artifacts in poorly sampled common-midpoint gathers. [yu2007wavelet] proposed a wavelet-Radon transform-based dealiasing algorithm that exploited the blurring of frequency information across wavelet scales to interpolate aliased signal components, exhibiting better performance than the Spitz method does. A new digital transform named seislet transform was proposed by [fomel2010seislet] to interpolate seismic data. It combined the wavelet-lifting scheme with local plane-wave destruction, providing a convenient orthogonal basis for denoising and interpolation. [naghizadeh2010beyond] proposed an interpolation method based on the curvelet transform for aliased seismic data. This study proves that spatially aliased data can be represented in the curvelet domain by two types of coefficients in coarser and finer scales. As a newly developed sparse transform at the time, the dreamlet transform was adopted by [wang2014dreamlet] to interpolate synthetic and field seismic data, exhibiting better performances in terms of sparsity and signal-to-noise ratio (SNR) than those of the curvelet transform-based method. Based on the good sparse representation ability of shearlet transform, [liu2018crossline] proposed a multi-component crossline seismic data reconstruction method based on sparse shearlet constraint inversion, achieving better results than traditional wavelet, curvelet, and shearlet methods in several extremely sparse sampling cases. Most of the above mathematical transform-based methods are based on fixed bases [liang2014seismic]. To learn basis functions adaptively, experts have developed dictionary learning-based interpolation methods. [liang2014seismic] attempted to restore decimated seismic data using data-driven tight frame (DDTF) first developed by [cai2014data]. The DDTF can adaptively learn from the data itself, and provides a sparser representation for the data in turn. [yu2015interpolation] extended the DDTF to high-dimensional versions and achieves the simultaneous denoising and interpolation of 3D and 5D seismic data. [wang2020fast] proposed an efficient dictionary learning method for high-dimensional seismic data reconstruction based on sequential generalized K-means model instead of the K-singular value decomposition (KSVD), achieving comparable interpolation performance to the traditional KSVD method at a very low computational cost. Nevertheless, these sparse-representation methods are hampered by inherent limitations, despite their excellent anti-aliasing and interpolation performance. First, the sparsity of data is an essential assumption. Second, the choice of parameters plays a crucial role. For instance, for the curvelet transform, inaccuracies in the estimation of the mask function can directly affect interpolation performance; for the seislet transform, inaccurate estimation of the local event slope tends to deteriorate the fidelity of seismic data reconstruction [wang2019deep].
@@
Generally, the successful applications of these conventional methods must satisfy a certain number of prior assumptions, including an accurate subsurface velocity model, the linearity of events, the low-rank structures of seismic data, and the sparsity of signals [jia2017what, wang2019deep, dong2025deep]. However, it is usually difficult to ensure that the understanding of the subsurface structure is consistent with real conditions [cheng2024meta], restricting the effectiveness and adaptability of these theory-driven interpolation approaches in handling several real and complex scenarios [jia2017what]. These theory-driven methods also involve complex and sophisticated operations for parameter adjustments, which extremely requires expert knowledge and hands-on experience. Additionally, the heavy computational cost that occasionally occurs is also a major hurdle, especially when handling with large-scale and high-dimensional seismic datasets [wang2019deep, dong2022seismic].
@@
In recent years, data-driven techniques have attracted much attention for their ability to learn from the data itself and process large amounts of data more efficiently [cheng2024meta]. As a representative data-driven technique, deep learning [DL; lecun2015deep] refers to a class of representation-learning methods that automatically mine useful features from raw data by stacking numerous nonlinear transformation layers. Each layer extracts increasingly abstract representations, enabling the system to learn complex functions directly from the input data without manual feature engineering [lecun2015deep, alzubaidi2021review, mousavi2022deep]. These characteristics enable DL-based methods perform well without prior assumptions and handcrafted parameters, so as to automatically process large-scale datasets. Following the pioneering efforts [mandelli2018seismic, mikhailiuk2018deep, wang2018seismic], researchers have increasingly explored DL-based approaches for seismic data interpolation, with convolutional neural networks (CNNs) being the primary focus in the early phase. Typical CNN-based architectures, such as U-Net [ronneberger2015unet, park2019reconstruction, park2021method, fang2021seismic], Res-Net [He_2016_CVPR, wang2018seismic, wang2019deep, liu2022aseismic], and generative adversarial network [GAN; goodfellow2014generative, siahkoohi2018seismic, oliveira2018interpolating], have shown great promise for the seismic data interpolation. Various improvements applied to CNNs have also contributed to further enhance the interpolation performance, including the attention-enhanced CNNs [yu2021attention], depthwise separable CNNs [jin2023depthwise], and multi-scale CNNs [cheng2023seismic, dong2024can, dong2024seismic]. As researchers have gained deeper insights into DL-based interpolation methods, the local perception has emerged as a major bottleneck of CNNs. This property leads to the neglect of global features and hinders the further interpolation performance enhancement using DL. To address this issue, experts have turned to Transformer [vaswani2017attention] owing to its strong ability to extract global contextual information using self-attention mechanism. [guo2023seismic] proposed to apply the window-based multi-head self-attention (MSA) into a multi-scale Transformer network. It can capture long-range dependencies to reconstruct more detailed information than CNNs do. Utilizing shifted window based MSA, [peng2023irregularly] trained an interpolation Swin-Transformer that combines Swin-Transformer modules and dense skip connections in a self-supervised manner, achieving higher accuracy and robustness than the convolutional autoencoder method. [cheng2024seismic] proposed a seismic interpolation Transformer network, which is an encoder-decoder structure, with Swin-Transformer blocks forming a U-shaped structure between the encoder and decoder. This approach can effectively reconstruct the consecutively missing traces in distributed acoustic sensing-vertical seismic profiling data. By integrating the global and local features, [gao2024swin] propose a Swin-Transformer convolutional residual network for the simultaneous denoising and interpolation of seismic data. This CNN-Transformer hybrid model has gain better visual performance and quantitative indices than conventional methods and CNN-based methods. A dense double branch attention Transformer (D2AT) proposed by [dong2025deep] can effectively reconstruct the consecutively missing traces. The core module of D2AT is a global feature unit composed of six Swin-Transformer blocks distributed in two scales, and it has shown great interpolation performance and generalization in both pre-stack and post-stack datasets.
@@
Although these DL-based methods have achieved superior performance over conventional methods, they primarily focus on two-dimensional (2D) interpolation and overlook the spatial structure correlation that could be exploited in a 3D interpolation workflow [wang2019deep, liu2022aseismic, dong2025deep]. However, 3D feature extraction operations usually introduce significantly greater computational burdens for DL-based interpolations. This phenomenon is particularly prominent when applying Transformer-based architectures, whose computational complexity scales quadratically with the size of input data [vyas2020fast]. Moreover, extending 2D networks to 3D versions will significantly increase the number of trainable parameters [ye20193d] . Several scholars have made attempts on 3D seismic data interpolation based on 3D CNNs in recent years [qian2021dtae, chen2023projection, saad2023unsupervised, wang2025self], but most of these methods suffer from heavy computational burdens. Obviously, this phenomenon will become more severe in 3D Transformers due to their quadratic computational complexity. To the best of our knowledge, no prior effort has attempted to apply Transformers to 3D seismic data interpolation at this stage, as training a 3D Transformer network on 3D seismic datasets demands prohibitively large computational resources. The application of Transformers to 3D seismic data interpolation remains an uncharted research area and holds significant potential for future investigations.
@@
A recent study [pan2022st] demonstrates a promising approach to address the above challenges. [pan2022st] proposed an image-to-video transfer learning (TL) strategy that can transfer an 2D image model to a 3D video model using parameter-efficient fine-tuning operations. Inspired by this cross-dimensional (i.e., 2D to 3D) TL strategy, we propose a lightweight 2.5-dimensional (2.5D) Transformer (T-2.5D) network to reconstruct 3D incomplete seismic data without using a large amount of 3D volumes to optimize a 3D Transformer network. The proposed T-2.5D is a hybrid of 2D and 3D modules, including four 2D Transformer encoders and four 3D seismic dimension adapters (SDAs). We design a 2D-to-3D cross-dimensional training workflow to optimize the T-2.5D. This workflow contains two stages: a 2D pre-training stage and a 3D fine-tuning stage. In the 2D pre-training stage, we use 2D training patches to just optimize the four 2D Transformer encoders without SDAs. In the 3D fine-tuning stage, we freeze the trainable parameters of the four 2D Transformer encoders and use limited 3D volumes to train the four 3D SDAs. This fine-tuning stage enables the T-2.5D to learn spatial correlation information of 3D seismic data. Overall, we use the ‘2D pre-training + 3D fine-tuning’ to replace the time-consuming and memory-intensive full 3D training, so as to significantly alleviate the computational cost of 3D Transformer-based interpolation. We investigate the interpolation performance of T-2.5D on several 3D seismic volumes. Experimental results show that the T-2.5D achieves comparable interpolation performance to 3D Transformer at a significantly lower computational cost. In other words, the proposed T-2.5D achieves a better trade-off between interpolation performance and computational cost.
The remainder of this paper is organized as follows. We begin with the theory of Transformer. Then, detailed descriptions of the proposed T-2.5D and the cross-dimensional TL are provided. Subsequently, we present the experimental results of several field examples. Finally, we draw a conclusion by summarizing our work.

### v2

@@
2.5D Transformer: An Efficient 3D Seismic Interpolation Method without Full 3D Training
@@
Machine Learning, Image processing, Neural networks, fuzzy logicTransformer has emerged as a powerful deep-learning technique for two-dimensional (2D) seismic data interpolation, owing to its global modeling ability. However, its core operation introduces heavy computational burden due to the quadratic complexity, hindering its further application to higher-dimensional data. To achieve Transformer-based three-dimensional (3D) seismic interpolation, we propose a 2.5-dimensional Transformer network (T-2.5D) that adopts a cross-dimensional transfer learning (TL) strategy, so as to adapt the 2D Transformer encoders to 3D seismic data. The proposed T-2.5D is mainly composed of 2D Transformer encoders and 3D seismic dimension adapters (SDAs). Each 3D SDA is placed before a Transformer encoder to learn spatial correlation information across seismic lines. The proposed cross-dimensional TL strategy comprises two stages: 2D pre-training and 3D fine-tuning. In the first stage, we optimize the 2D Transformer encoders using a large amount of 2D data patches. In the second stage, we freeze the 2D Transformer encoders and fine-tune the 3D SDAs using limited 3D data volumes. Extensive experiments on multiple datasets are conducted to assess the effectiveness and efficiency of T-2.5D. Experimental results demonstrate that the proposed method achieves comparable performance to that of full 3D Transformer at a significantly low cost.
1 Introduction
@@
Interpolation operation is a crucial step for improving the quality of seismic data, especially when facing geometries with large spatial sampling intervals, and recovering missing parts caused by topographical constraints, limited budgets, or receiver malfunctions. By reconstructing the missing information and compensating the sparse geometries, effective interpolations can enhance the completeness and continuity of events, facilitating the subsequent inversion and high-resolution imaging (Chen et al., 2024; Cheng et al., 2025).
@@
To acquire the desired dense and complete seismic data, geophysical scholars have proposed various conventional interpolation methods over the past few decades (Spitz, 1991; Trad et al., 2002; Oropeza and Sacchi, 2011) . These methods are theory-driven, relying on mathematical models or analytical formulations to describe the underlying physical processes. Generally, they can be divided into four categories: the wave-equation-based, prediction filtering, low-rank, and sparse-representation methods.
@@
The wave-equation-based methods utilize subsurface velocity models to simulate the propagation of seismic wavefields, so as to reconstruct the incomplete data (Wang et al., 2019). Ronen (1987) proposed a trace-interpolation method based on the wave equation and prior assumption of a smooth spatial spectrum, and experimental results have demonstrated its effectiveness on both synthetic and field data. The offset-continuation differential equation-based method proposed by Fomel (2003) showed good interpolation performance even in several structurally complex situations. However, the wave-equation based methods often require a relatively accurate velocity model as an important prerequisite, which is often unavailable in numerous real-world problems. Moreover, the heavy computation burden seriously limits their further applications.
@@
The prediction filtering methods exploit the correlation of events in different domains to perform the interpolation operation. When applying these methods, the interpolation task is taken as a least-squares linear inverse problem, and interpolated results are generated by minimizing the misfit between the predicted and input data (Liu et al., 2022b). Spitz (1991) proposed a multichannel and model-free interpolation method for missing traces in f-x domain, exhibiting good interpolation performance in both pre- and post-stack examples. To further improve the interpolation performance, Naghizadeh and Sacchi (2007) proposed a multistep autoregressive algorithm-based prediction-error method combined with Fourier-based methods to reconstruct seismic data from low to all frequencies. An important assumption of these prediction filtering methods is that events are required to be linear. However, this is usually not satisfied in real conditions, resulting in serious interpolation performance degradation.
@@
The low-rank methods assume that the complete seismic data in a specific arrangement is of low rank at a given frequency component (Trickett et al., 2010; Gao et al., 2013), and missing traces will significantly increase its rank. Therefore, the interpolation tasks can be accomplished by reducing the rank to an optimal one. Typical low-rank methods include singular spectrum analysis (SSA; Vautard and Ghil, 1989; Oropeza and Sacchi, 2011; Carozzi and Sacchi, 2021), matrix completion (Ma, 2013; Yang et al., 2013; Kumar et al., 2015), principal component analysis (Wold et al., 1987; Huang et al., 2016; Wu et al., 2023), and Cadzow filtering (Cadzow, 2002; Gao et al., 2013; Naghizadeh and Sacchi, 2013; Huang et al., 2020). Oropeza and Sacchi (2011) proposed a simultaneous denoising and reconstruction method for seismic data based on multichannel SSA (MSSA), which resembles seismic data interpolation with the method of projection onto convex sets (POCS). Moreover, a randomized singular value decomposition is adopted to accelerate its rank reduction stage. Ma (2013) enhanced the matrix completion using a designed texture-patch transformation, exhibiting superior performance to traditional POCS method. However, in these low-rank methods, the determination of optimal rank is an intricate problem, and an improper one will have a negative impact on the interpolated results (Ma, 2013; Cheng et al., 2023).
@@
The sparse-representation methods interpolate seismic records by representing them into sparse domains. The useful signals and the missing data are treated as large- and small-amplitude coefficients, respectively. Preserving the large-amplitude coefficients and eliminating the small-amplitude ones helps to extract the useful signal components from the incomplete data (Chen et al., 2019). The missing signal components are then recovered through an inverse sparse transform. These methods can be mainly categorized into mathematical transform-based (Trad et al., 2002; Yu et al., 2007; Fomel and Liu, 2010; Naghizadeh and Sacchi, 2010) and dictionary learning-based (Liang et al., 2014; Yu et al., 2015; Wang et al., 2020) methods. Trad et al. (2002) proposed a high-resolution time-variant Radon transform (RT)-based interpolation method. Hyperbolic and elliptical RTs are implemented to perform accurate interpolation and attenuate sampling artifacts in poorly sampled common-midpoint gathers. Based on the good sparse representation ability of shearlet transform, Liu et al. (2018) proposed a multi-component crossline seismic data reconstruction method based on sparse shearlet constraint inversion, achieving better results than traditional wavelet, curvelet, and shearlet methods in several extremely sparse sampling cases. However, most of the above mathematical transform-based methods are based on fixed bases (Liang et al., 2014). To learn basis functions adaptively, experts have developed dictionary learning-based interpolation methods. Liang et al. (2014) attempted to restore decimated seismic data using data-driven tight frame (DDTF) first developed by Cai et al. (2014). The DDTF can adaptively learn from the data itself, and provides a sparser representation for the data in turn. Yu et al. (2015) extended the DDTF to high-dimensional versions and achieves the simultaneous denoising and interpolation of 3D and 5D seismic data. Nevertheless, these sparse-representation methods are hampered by inherent limitations, despite their excellent anti-aliasing and interpolation performance. First, the sparsity of data is an essential assumption. Second, the choice of parameters plays a crucial role. For instance, for the curvelet transform, inaccuracies in the estimation of the mask function can directly affect interpolation performance; for the seislet transform, inaccurate estimation of the local event slope tends to deteriorate the fidelity of seismic data reconstruction. (Wang et al., 2019).
@@
Generally, the successful applications of these conventional methods must satisfy a certain number of prior assumptions, including an accurate subsurface velocity model, the linearity of events, the low-rank structures of seismic data, and the sparsity of signals (Jia and Ma, 2017; Wang et al., 2019; Dong et al., 2025). However, it is usually difficult to ensure that the understanding of the subsurface structure is consistent with real conditions (Cheng et al., 2024b), restricting the effectiveness and adaptability of these theory-driven interpolation approaches in handling several real and complex scenarios (Jia and Ma, 2017). These theory-driven methods also involve complex and sophisticated operations for parameter adjustments, which extremely requires expert knowledge and hands-on experience. Additionally, the heavy computational cost that occasionally occurs is also a major hurdle, especially when handling with large-scale and high-dimensional seismic datasets (Wang et al., 2019; Dong et al., 2022).
@@
In recent years, deep learning (DL; LeCun et al., 2015) have attracted much attention. This kind of data-driven techniques has the ability to learn from data itself and process large amounts of data more efficiently (Cheng et al., 2024b). These characteristics enable DL-based methods perform well without prior assumptions and handcrafted parameters, so as to automatically process large-scale datasets. At the early stage, convolutional neural networks (CNNs) have been the primary focus. Typical CNN-based architectures, such as U-Net (Ronneberger et al., 2015; Park et al., 2019, 2021; Fang et al., 2021), Res-Net (He et al., 2016; Wang et al., 2018, 2019; Liu et al., 2022a), and generative adversarial network (GAN; Goodfellow et al., 2014; Siahkoohi et al., 2018; Oliveira et al., 2018), have shown great promise for the seismic data interpolation. Various improvements applied to CNNs have also contributed to further enhance the interpolation performance, including the attention-enhanced CNNs (Yu and Wu, 2021), depthwise separable CNNs (Jin et al., 2023), and multi-scale CNNs (Cheng et al., 2023; Dong et al., 2024a, b). As researchers have gained deeper insights into DL-based interpolation methods, the local perception has emerged as a major bottleneck of CNNs, which hinders the further enhancement of the interpolation performance using DL. To address this issue, experts have turned to Transformer (Vaswani et al., 2017) owing to its strong ability to extract global contextual information using self-attention mechanism. Cheng et al. (2024a) proposed a seismic interpolation Transformer network, which is an encoder-decoder structure, with Swin-Transformer blocks forming a U-shaped structure between the encoder and decoder. This approach can effectively reconstruct the consecutively missing traces in distributed acoustic sensing-vertical seismic profiling data. By integrating the global and local features, Gao et al. (2024) proposed a Swin-Transformer convolutional residual network for the simultaneous denoising and interpolation of seismic data, gaining better visual performance and quantitative indices than conventional methods and CNN-based methods. A dense double branch attention Transformer (D2AT) proposed by Dong et al. (2025) can effectively reconstruct the consecutively missing traces. The core module of D2AT is a global feature unit composed of six Swin-Transformer blocks distributed in two scales, and this method has shown great interpolation performance and generalization in both pre-stack and post-stack datasets.
@@
Although these DL-based methods have achieved superior performance over conventional methods, they primarily focus on two-dimensional (2D) interpolation and overlook the spatial structure correlation that could be exploited in a 3D interpolation workflow (Wang et al., 2019; Liu et al., 2022a; Dong et al., 2025). However, 3D feature extraction operations usually introduce significantly greater computational burdens for DL-based interpolations. This phenomenon is particularly prominent when applying Transformer-based architectures, whose computational complexity scales quadratically with the size of input data (Vyas et al., 2020). Moreover, extending 2D networks to 3D versions will significantly increase the number of trainable parameters (Ye et al., 2019). Several scholars have made attempts on 3D seismic data interpolation based on 3D CNNs in recent years (Qian et al., 2021; Chen et al., 2023; Saad et al., 2023; Wang et al., 2025), but most of these methods suffer from heavy computational burdens. Obviously, this phenomenon will become more severe in 3D Transformers due to their quadratic computational complexity. To the best of our knowledge, no prior effort has attempted to apply Transformers to 3D seismic data interpolation at this stage, as training a 3D Transformer network on 3D seismic datasets demands prohibitively large computational resources.
@@
A recent study (Pan et al., 2022) demonstrates a promising approach to address the above challenges. Pan et al. (2022) proposed an image-to-video transfer learning (TL) strategy that can transfer an 2D image model to a 3D video model using parameter-efficient fine-tuning operations. Inspired by this, we propose a lightweight 2.5-dimensional (2.5D) Transformer (T-2.5D) network to reconstruct 3D incomplete seismic data without using a large amount of 3D volumes to optimize a 3D Transformer network. The proposed T-2.5D is a hybrid of 2D and 3D modules, including four 2D Transformer encoders and four 3D seismic dimension adapters (SDAs). We design a 2D-to-3D cross-dimensional TL workflow to optimize the T-2.5D. This workflow contains two stages: a 2D pre-training stage and a 3D fine-tuning stage. In the 2D pre-training stage, we use 2D training patches to just optimize the four 2D Transformer encoders without SDAs. In the 3D fine-tuning stage, we freeze the trainable parameters of the four 2D Transformer encoders and use limited 3D volumes to train the four 3D SDAs. This fine-tuning stage enables the T-2.5D to learn spatial correlation information of 3D seismic data. Overall, we use the ‘2D pre-training + 3D fine-tuning’ to replace the time-consuming and memory-intensive full 3D training, so as to significantly alleviate the computational cost of 3D Transformer-based interpolation. We investigate the interpolation performance of T-2.5D on several 3D seismic volumes. Experimental results show that the T-2.5D achieves comparable interpolation performance to 3D Transformer at a significantly lower computational cost. In other words, the proposed T-2.5D achieves a better trade-off between interpolation performance and computational cost.